OpenStudy (anonymous):

Determine the conversion factor for a conversion of m/s2 to km/h2? 1 kg = 2.205 lb, 1 in. = 2.54 cm, 12 in. = 1 ft 1 m/s^2 =______ km/h^2 ??

OpenStudy (anonymous):

@phi @jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: 1 km = 1000 m and 3600 s = 1 hr (since 60 s = 1 min, 60 min = 1 hr, so 60*60 = 3600 s = 1 hr)

OpenStudy (anonymous):

but it's h^2 and s^2? would it be 3600*1000?

jimthompson5910 (jim_thompson5910):

I want to say you square both sides to get 3600 s = 1 hr (3600 s)^2 = (1 hr)^2 12,960,000 s^2 = 1 hr^2 but that sounds awfully large

OpenStudy (anonymous):

that's not it according to the problem

OpenStudy (anonymous):

it's km over h^2 so would you divide?

OpenStudy (anonymous):

yep that's it!

OpenStudy (anonymous):

thanks though!

jimthompson5910 (jim_thompson5910):

alright glad you figured it out